# A mesh-free hybrid Chebyshev-Tucker tensor format with applications to multi-particle modelling

### v1

@@ 4 Applications to multi-particle modelling
In this section, we consider the calculation of a weighted sum of interaction potential in a $N$-particle system in $\mathbb{R}^{3}$:
$P(x)=\sum_{v=1}^{N}z_{v}p(\|x-x_{v}\|),\quad z_{v}\in\mathbb{R}\text{ and }x_{v},x\in[-1,1]^{d},$ (4.1)
[…]
@@
In what follows, we consider the case of Newton generating kernel $p(\|x\|)=1/\|x\|$ for $d=3$, arising in electronic structure calculations and in bio-molecular modeling. We first recall the results on a Sinc-quadrature based constructive CP tensor representation of a single Newton kernel discretized on a 3D Cartesian grid [21, 9], thereby providing the CP tensor input for our mesh-free ChebTuck approximation Algorithm 4. We then introduce the idea of range-separation for efficient numerical treatment of the single Newton kernel [4]. Finally, we discuss how the total potential Eq. 4.1 can be constructed from the CP tensor representation of the single Newton kernel, for both bio-molecular modeling and lattice-structured compounds with vacancies, for which cases our ChebTuck approximation can be gainfully applied.
@@ 4.1.1 A low-rank CP tensor representation of the Newton kernel
We recall the results on a constructive low-rank representation of the Newton kernel developed in [21, 9] based on sinc-quadrature. Let ${\bf P}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}$ be the projection-collocation tensor of the Newton kernel $p(x)=1/\|x\|$ discretized on a $n_{1}\times n_{2}\times n_{3}$ 3D uniform Cartesian grid in $[-1,1]^{3}$, where the entries of ${\bf P}=\{{p}_{{\bf i}}\}$ are defined by
${p}_{{\bf i}}=\int_{[-1,1]^{3}}\dfrac{\psi_{{\bf i}}(x)}{\|x\|}\operatorname{d}x=\dfrac{1}{h^{3}}\int_{t_{i_{1}-1}^{(1)}}^{t_{i_{1}}^{(1)}}\int_{t_{i_{2}-1}^{(2)}}^{t_{i_{2}}^{(2)}}\int_{t_{i_{3}-1}^{(3)}}^{t_{i_{3}}^{(3)}}\dfrac{1}{\|x\|}\operatorname{d}x_{1}\operatorname{d}x_{2}\operatorname{d}x_{3},$ (4.2)
@@
with the multi-index ${\bf i}=(i_{1},i_{2},i_{3})\in\{1,\cdots,n\}^{3}$, where $t_{i_{\ell}}^{(\ell)}=-1+(i_{\ell}-1)h_{\ell}$ is the grid points in each dimension and $h_{\ell}=2/(n_{\ell}-1)$ is the grid size. The basis function $\psi_{{\bf i}}$ is a product of piecewise constant functions in each dimension, i.e.,
$\psi_{{\bf i}}(x)=\psi_{i_{1}}(x_{1})\psi_{i_{2}}(x_{2})\psi_{i_{3}}(x_{3}),$ (4.3)
where $\psi_{i_{\ell}}(x_{\ell})=\chi_{[t_{i_{\ell-1}}^{(\ell)},t_{i_{\ell}}^{(\ell)})}(x_{\ell})/h_{\ell}$ is the piecewise constant function that is $1/h_{\ell}$ on the interval $[t_{i_{\ell-1}}^{(\ell)},t_{i_{\ell}}^{(\ell)})$ and 00 elsewhere. To ease the notations, we assume that $n_{1}=n_{2}=n_{3}=n$ in the following and therefore we use $h$ and $t_{i_{\ell}}$ to denote the grid size and grid points in all dimensions respectively.
We refer to [21, 9] for the detailed construction of tensor ${\bf P}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}$, its sinc-based CP approximation and the corresponding error analysis. A numerically optimized procedure for rank minimization of the sinc-based CP approximation of ${\bf P}$ is then applied [9], leading to the final rank-$R$ CP decomposition ${\bf P}_{R}$ of the collocation tensor ${\bf P}$
${\bf P}\approx{\bf P}_{R}=\sum_{k=1}^{R}{\bf p}_{k}^{(1)}\otimes{\bf p}_{k}^{(2)}\otimes{\bf p}_{k}^{(3)}$ (4.4)
where ${\bf p}_{k}^{(\ell)}$ are canonical vectors. We remark here that the canonical vectors in each dimension are the same due to the radial symmetry of the Newton kernel, i.e., ${\bf p}_{k}^{(1)}={\bf p}_{k}^{(2)}={\bf p}_{k}^{(3)}={\bf p}_{k}$. Note that here the rank parameter $R$ is adapted to the chosen approximation error and to the corresponding grid size $n$, and it scales only logarithmically in both parameters. Similarly CP tensor decomposition can be constructed for a wide class of analytic kernels, for example, for the discretized Slater potential, $\mathrm{e}^{-\lambda\|x\|}$, $x\in\mathbb{R}^{3}$.
Figure 4.1: The canonical vectors ${\bf p}_{k}^{(1)}$ for $k=1,\cdots,R$, $R=26$, along the $x$-axis for the rank-$R$ CP tensor approximation of collocation tensor of size ${\bf P}_{R}\in\mathbb{R}^{n\times n\times n}$ with $n=256$.
@@
In Fig. 4.1, we illustrate the CP approximation of the collocation tensor on the example of the Newton kernel discretized on a $n\times n\times n$ 3D Cartesian grid for $n=256$, presented for example in [30, 26]. In this case, $R=26$ gives a relative error $10^{-7}$. We plot the canonical vectors ${\bf p}_{k}^{(1)}$ for $k=1,\cdots,26$ along the $x$-axis. We can see that there are canonical vectors localized around the singularity at the origin as well as slowly decaying ones, which clearly corresponds to the long-range and short-range behavior of the Newton kernel. We remark here that the number of vectors $R$ depends logarithmically on the size of $n\times n\times n$ 3D Cartesian grid and the parameter $\varepsilon>0$ controlling the accuracy of sinc-approximation. For increasing grid size $n$ the number of vectors with localized support around the singularity at the origin becomes larger. Choosing smaller $\varepsilon$-parameter also affects the slow logarithmic increase of the number of required canonical vectors for the tensor representation of the given kernel.
@@
This low-rank CP tensor representation of the Newton kernel has been successfully used by two of coauthors in high accuracy electronic structure calculations [31, 26]. In particular, it was applied for tensor based computation of the Hartree potential (3D convolution of the electron density with the Newton kernel), the nuclear potential operators, and the two-electron integrals tensor, see [26] and references therein. All computations have been performed using huge 3D $n\times n\times n$ Cartesian grids with $n$ of the order of $10^{4}\sim 10^{5}$, providing accuracy close to the results from the analytically based quantum chemical program packages.
4.1.2 Why the Chebyshev polynomial interpolation fails to approximate ${\bf P}_{R}$ without range-separation
@@
In this subsection, we consider the ChebTuck approximation Algorithm 4 applied to the CP tensor ${\bf P}_{R}$ of the singular Newton kernel and discuss why it may fail to give satisfactory results, which necessitates the use of the range-separated (RS) tensor format [4].
@@
$\mathrm{err}:=\max_{i_{1},i_{2}}\left|{\bf P}_{R}(i_{1},i_{2},n/2)-\hat{f}_{\bf m}(t_{i_{1}},t_{i_{2}},t_{n/2})\right|/\max_{i_{1},i_{2}}|{\bf P}_{R}(i_{1},i_{2},n/2)|,$ (4.5)
as shown in Table 4.1. We can see that in this case, only when $m\gg n$ the ChebTuck approximation can give satisfactory results, which is not practical and does not reduce the storage cost.
@@
256 0.41 0.07 $7.7\cdot 10^{-3}$ $6.9\cdot 10^{-4}$ $2.0\cdot 10^{-4}$ $8.5\cdot 10^{-6}$ $1.6\cdot 10^{-6}$ $3.5\cdot 10^{-7}$
512 0.71 0.41 0.07 $7.7\cdot 10^{-3}$ $6.9\cdot 10^{-4}$ $2.0\cdot 10^{-4}$ $8.5\cdot 10^{-6}$ $1.6\cdot 10^{-6}$
1024 0.92 0.71 0.41 0.07 $7.7\cdot 10^{-3}$ $6.9\cdot 10^{-4}$ $2.0\cdot 10^{-4}$ $8.5\cdot 10^{-6}$
2048 1.06 0.92 0.71 0.41 0.07 $7.7\cdot 10^{-3}$ $6.9\cdot 10^{-4}$ $2.0\cdot 10^{-4}$
4096 1.16 1.06 0.92 0.71 0.41 0.07 $7.7\cdot 10^{-3}$ $6.9\cdot 10^{-4}$
Table 4.1: Relative $\ell_{\infty}$ error in the middle slice of the ChebTuck approximation of the single Newton kernel as in Eq. 4.5 for different grid size $n$ and Chebyshev degree $m$.
@@
$\mathrm{err}:=\max_{i}\left|{\bf p}_{k}(i)-g_{k}(t_{i})\right|/\max_{i}|{\bf p}_{k}(i)|,$ (4.6)
for different Chebyshev degree $m$. The results are shown in Fig. 4.2. We can see that canonical vectors representing the long-range behavior of the Newton kernel, i.e., ${\bf p}_{1},\cdots,{\bf p}_{15}$ can be well approximated by low-degree Chebyshev interpolants (left), while for short-range canonical vectors, the convergence is much slower (middle). This is somewhat expected since the short-range canonical vectors have singularities at the origin and require high-degree polynomials to capture the behavior. In Fig. 4.2 right, we plot the Chebyshev approximation error of different canonical vectors for a fixed Chebyshev degree. We can see given this fixed degree $m=129$, the error of the short-range functions is much larger than the long-range functions with a jump around $k=15$. The same behavior can also be seen from Fig. 4.3, in which the absolute values of the Chebyshev coefficients of the interpolant of a long-range vector ${\bf p}_{10}$ and a short-range vector ${\bf p}_{30}$ are shown. We can see that for the long-range vector, the Chebyshev coefficients decay exponentially while for the short-range vector, the coefficients do not decay very much.
@@
Figure 4.2: Relative $\ell_{\infty}$ error Eq. 4.6 of the Chebyshev interpolation of the canonical vectors v.s. the degree of the Chebyshev interpolant. Each line represents the error of the Chebyshev interpolant of a canonical vector ${\bf p}_{k}$ for $k=1,\cdots,15$ (left) and $k=16,\cdots,39$ (middle). Right: Relative $\ell_{\infty}$ error of the degree $m=129$ Chebyshev interpolation of the canonical vectors v.s. the index $k$ of the univariate function.
Figure 4.3: Absolute values of the Chebyshev coefficients of the degree-129 Chebyshev interpolant of ${\bf p}_{10}$ (left) and ${\bf p}_{30}$ (right). Only coefficients of even degree are shown, i.e., ${\bf c}_{k}(0),{\bf c}_{k}(2),\cdots,{\bf c}_{k}(128)$ since the Newton kernel is even and ${\bf c}_{k}(1),{\bf c}_{k}(3),\cdots$ are zero.
@@ 4.1.3 Range-separation of the CP tensor ${\bf P}_{R}$
For given an appropriately chosen range-separation parameter $R_{s}\in[1,R]\cap\mathbb{Z}$ (see [4] for details of the strategies for choosing $R_{s}$), we further use the range-separated representation of the CP tensor ${\bf P}_{R}$ constructed in [4] in the form,
${\bf P}_{R}={\bf P}_{R_{s}}+{\bf P}_{R_{l}},$ (4.7)
where
${\bf P}_{R_{s}}=\sum_{k=1}^{R_{s}}{\bf p}_{k}\otimes{\bf p}_{k}\otimes{\bf p}_{k},\quad{\bf P}_{R_{l}}=\sum_{k=R_{s}+1}^{R}{\bf p}_{k}\otimes{\bf p}_{k}\otimes{\bf p}_{k}.$ (4.8)
Here Eq. 4.8 represents the sums over the sets of indexes for the long- and short-range canonical vectors in the range-separation splitting of the Newton kernel. Then the short-range part ${\bf P}_{R_{s}}$ is a highly localized tensor which has very small support and can thus be stored sparsely (see [4] for details), while the ChebTuck approximation can be gainfully applied to the long-range part ${\bf P}_{R_{l}}$.
We apply Algorithm 4 to the long-range part ${\bf P}_{R_{l}}$ and measure the errors of the approximation at the middle slice as in Eq. 4.5 for different grid size $n$ and Chebyshev degree $m$, which are shown in Table 4.2. As opposed to the Table 4.1, we can see that the ChebTuck approximation of the long-range part ${\bf P}_{R_{l}}$ can give satisfactory results already for reasonable Chebyshev degrees $m$. We also show the optimal Tucker ranks, i.e., size of the core tensor $\bm{\beta}$ of the ChebTuck approximation of the long-range part ${\bf P}_{R_{l}}$ computed by RHOSVD with relative error $10^{-7}$ for different grid size $n$ and Chebyshev degree $m$ in Table 4.3. For fixed Chebyshev degree, the Tucker rank of the long-range part only increases logarithmically with the grid size $n$. This indicates that the long-range part can be well approximated by a ChebTuck format with a small number of parameters.
$n$ $m$ 129 257 513 1025 2049 4097
256 $8.4\cdot 10^{-8}$ $7.4\cdot 10^{-8}$ $7.3\cdot 10^{-8}$ $4.9\cdot 10^{-9}$ $1.6\cdot 10^{-9}$ $1.8\cdot 10^{-10}$
512 $5.2\cdot 10^{-8}$ $5.2\cdot 10^{-8}$ $4.6\cdot 10^{-8}$ $4.6\cdot 10^{-8}$ $3.0\cdot 10^{-9}$ $9.9\cdot 10^{-10}$
1024 $9.4\cdot 10^{-9}$ $9.4\cdot 10^{-9}$ $9.4\cdot 10^{-9}$ $8.7\cdot 10^{-9}$ $8.6\cdot 10^{-9}$ $5.7\cdot 10^{-10}$
2048 $3.6\cdot 10^{-9}$ $1.7\cdot 10^{-9}$ $1.7\cdot 10^{-9}$ $1.7\cdot 10^{-9}$ $1.6\cdot 10^{-9}$ $1.6\cdot 10^{-9}$
4096 $1.1\cdot 10^{-4}$ $7.5\cdot 10^{-10}$ $7.4\cdot 10^{-10}$ $7.4\cdot 10^{-10}$ $7.4\cdot 10^{-10}$ $7.1\cdot 10^{-10}$
Table 4.2: Relative $\ell_{\infty}$ error as Eq. 4.5 in the middle slice of the ChebTuck approximation of the long-range part ${\bf P}_{R_{l}}$ for different grid size $n$ and degree $m$.
$n$ $m$ 129 257 513 1025 2049 4097
256 9 9 9 9 9 9
512 11 11 11 11 11 11
1024 12 12 12 12 12 12
2048 13 13 13 13 13 13
4096 15 15 15 15 15 15
Table 4.3: Tucker rank $\max_{\ell=1,2,3}{r_{i_{\ell}}}$ of the ChebTuck format with RHOSVD truncation tolerance $10^{-7}$ for the long-range part of the Newton kernel for different grid size $n$ and Chebyshev degree $m$.

### v2

@@ 4 Applications to multi-particle modeling
In this section, we apply our hybrid tensor format to the challenging problem of approximating multi-particle interaction potentials. We consider the total potential generated by an $N$-particle system in $\mathbb{R}^{3}$, given by a weighted sum
(33) $P(x)={\sum}_{v=1}^{N}z_{v}p(\|x-x_{v}\|),\quad z_{v}\in\mathbb{R}\text{ and }x_{v},x\in[-1,1]^{d},$
where $x_{v}$ is the position of the $v$-th particle, $z_{v}$ is its weight and $p(\|x\|)$ is a radially symmetric generating kernel. These kernels are typically slowly decaying and singular at the origin. Examples include the Newton (Coulomb) $1/\|x\|$, Slater $\mathrm{e}^{-\lambda\|x\|}$, and Yukawa $\mathrm{e}^{-\lambda\|x\|}/\|x\|$ potentials, as well as Matérn covariance functions arising in various applications. Such practically interesting kernels have local singularities or cusps at the particle centers, and at the same time possess a non-local behavior that is responsible for the global interaction between all the particles in the system. Due to this short-long range behavior of the total collective potential $P(x)$ it does not allow for a low-rank tensor approximation in traditional tensor formats with satisfactory accuracy. This difficulty can be resolved by applying the RS tensor format [6].
@@
In what follows, we will focus on the ubiquitous Newton kernel $p(\|x\|)=1/\|x\|$ for $d=3$, arising in electronic structure calculations and in bio-molecular modeling. We first recall shortly the results on a sinc-quadrature based constructive CP tensor representation of a single Newton kernel discretized on a 3D Cartesian grid [28, 11], thereby providing the CP tensor input for our mesh-free ChebTuck approximation Algorithm 3. We then introduce the idea of range-separation for efficient numerical treatment of the single Newton kernel [6]. Finally, we discuss how the total potential Eq. 33 can be constructed from the CP tensor representation of the single Newton kernel, for both bio-molecular modeling and lattice-structured compounds with vacancies, for which cases our ChebTuck approximation can be gainfully applied.
@@ 4.1.1 A low-rank CP tensor representation of the Newton kernel
We first recall the constructive low-rank representation of the Newton kernel based on sinc-quadrature, as developed in [28, 11], and refer to these works for the detailed error analysis and numerically optimized procedure for rank minimization.
@@
Let ${\bf P}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}$ be the projection-collocation tensor of the Newton kernel $p(x)=1/\|x\|$ discretized on a $n_{1}\times n_{2}\times n_{3}$ 3D uniform Cartesian grid in $1,1]^{3}$, where the entries are defined by
(34) ${\bf P}_{{\bf i}}=\int_{[-1,1]^{3}}\dfrac{\psi_{{\bf i}}(x)}{\|x\|}\operatorname{d}x=\dfrac{1}{h_{1}h_{2}h_{3}}\int_{t_{i_{1}-1}^{(1)}}^{t_{i_{1}}^{(1)}}\int_{t_{i_{2}-1}^{(2)}}^{t_{i_{2}}^{(2)}}\int_{t_{i_{3}-1}^{(3)}}^{t_{i_{3}}^{(3)}}\dfrac{1}{\|x\|}\operatorname{d}x_{1}\operatorname{d}x_{2}\operatorname{d}x_{3},$
@@
with the multi-index ${\bf i}=(i_{1},i_{2},i_{3})\in\{1,\cdots,n\}^{3}$, where $t_{i_{\ell}}^{(\ell)}=-1+(i_{\ell}-1)h_{\ell}$ are the grid points in each dimension and $h_{\ell}=2/(n_{\ell}-1)$ is the grid size. The basis function $\psi_{{\bf i}}$ is a product of piecewise constant functions in each dimension $\psi_{{\bf i}}(x)=\psi_{i_{1}}(x_{1})\psi_{i_{2}}(x_{2})\psi_{i_{3}}(x_{3})$ where $\psi_{i_{\ell}}(x_{\ell})=\chi_{[t_{i_{\ell-1}}^{(\ell)},t_{i_{\ell}}^{(\ell)})}(x_{\ell})/h_{\ell}$. To ease notation, we assume that $n_{1}=n_{2}=n_{3}=n$ in the following, and therefore use $h$ and $t_{i_{\ell}}$ to denote the grid size and grid points in all dimensions respectively. A numerically optimized procedure for rank minimization of the sinc-based CP approximation of $\bf P$} is then applied, leading to the final rank-$R$ CP format ${\bf P}_{R}$ of the collocation tensor $\bf P$}
(35) ${\bf P}\approx{\bf P}_{R}={\sum}_{k=1}^{R}{\bf p}_{k}^{(1)}\otimes{\bf p}_{k}^{(2)}\otimes{\bf p}_{k}^{(3)}$
@@
where ${\bf p}_{k}^{(\ell)}$ are canonical vectors. We remark here that the canonical vectors in each dimension are the same due to the radial symmetry of the Newton kernel, i.e., ${\bf p}_{k}^{(1)}={\bf p}_{k}^{(2)}={\bf p}_{k}^{(3)}={\bf p}_{k}$. Note that here, the rank $R$ is adapted to the chosen approximation error and to the corresponding grid size $n$, and scales only logarithmically in both parameters. Similarly, the CP tensor decomposition can be constructed for a wide class of analytic kernels as mentioned in [28, 11]. In Fig. 1 left, we plot the canonical vectors ${\bf p}_{k}^{(1)}$ for $k=1,\ldots,R$ of the CP approximation ${\bf P}_{R}$ of $\bf P$} discretized on a $n\times n\times n$ Cartesian grid for $n=256$, presented for example in [37, 34]. In this case, $R=26$ gives a relative error in the order of $10^{-7}$. We can see that there are canonical vectors localized around the singularity at the origin as well as slowly decaying ones, which clearly corresponds to the long-range and short-range behavior of the Newton kernel. We remark here that the number of vectors $R$ depends logarithmically on the size of the $n\times n\times n$ 3D Cartesian grid and the parameter $\varepsilon>0$ controlling the accuracy of the sinc-approximation. For increasing grid size $n$, the number of vectors with localized support around the singularity at the origin becomes larger.
Figure 1: Left: the canonical vectors ${\bf p}_{k}^{(1)}$ for $k=1,\ldots,R$, $R=26$, along the $x$-axis for the rank-$R$ CP tensor approximation of collocation tensor of size ${\bf P}_{R}\in\mathbb{R}^{n\times n\times n}$ with $n=256$. Right: canonical vectors ${\bf a}_{k}^{(1)}$ of the long-range part in the total potential ${\bf A}_{R_{l}}\in\mathbb{R}^{n\times n\times n}$ with $N=500$ particles and grid size $n=256$ for $k=1,\ldots,8$.
@@
This low-rank CP tensor representation of the Newton kernel has been successfully used by two of the coauthors in high accuracy electronic structure calculations [39, 34]. In particular, it was applied for tensor based computation of the Hartree potential (3D convolution of the electron density with the Newton kernel), the nuclear potential operators, and the two-electron integral tensor, see [34] and the references therein. All computations have been performed using huge 3D $n\times n\times n$ Cartesian grids with $n$ of the order of $10^{4}\sim 10^{5}$, providing accuracy close to the results from the analytically based quantum chemical packages.
4.1.2 Why Chebyshev interpolation fails to approximate ${\bf P}_{R}$ without range-separation
@@
In this subsection, we consider the ChebTuck approximation Algorithm 3 applied to the CP tensor ${\bf P}_{R}$ of the singular Newton kernel and discuss why it fails to give satisfactory results, which necessitates the use of the range-separated (RS) tensor format [6].
@@
(36) $\mathrm{err}:=\max_{i_{1},i_{2}}\left|{\bf P}_{R}(i_{1},i_{2},n/2)-\hat{f}_{\bf m}(t_{i_{1}},t_{i_{2}},t_{n/2})\right|/\max_{i_{1},i_{2}}|{\bf P}_{R}(i_{1},i_{2},n/2)|,$
as shown in Table 1. We can see that in this case, only when $m\gg n$ the ChebTuck approximation provides satisfactory results, which is not practical and does not reduce the storage cost.
@@
256 0.41 0.07 7.7e-3 6.9e-4 2.0e-4 8.5e-6 1.6e-6 3.5e-7
512 0.71 0.41 0.07 7.7e-3 6.9e-4 2.0e-4 8.5e-6 1.6e-6
1024 0.92 0.71 0.41 0.07 7.7e-3 6.9e-4 2.0e-4 8.5e-6
2048 1.06 0.92 0.71 0.41 0.07 7.7e-3 6.9e-4 2.0e-4
4096 1.16 1.06 0.92 0.71 0.41 0.07 7.7e-3 6.9e-4
Table 1: Relative $\ell_{\infty}$ error in the middle slice of the ChebTuck approximation of the single Newton kernel as in Eq. 36 for different grid size $n$ and Chebyshev degree $m$.
@@
(37) $\mathrm{err}:=\max_{i}\left|{\bf p}_{k}(i)-g_{k}(t_{i})\right|/\max_{i}|{\bf p}_{k}(i)|,$
for different Chebyshev degree $m$. The results are shown in Fig. 2. We can see that canonical vectors representing the long-range behavior of the Newton kernel, i.e., ${\bf p}_{1},\ldots,{\bf p}_{15}$ can be well approximated by low-degree Chebyshev interpolants (left), while for short-range canonical vectors, the convergence is much slower (middle). This is as expected since the short-range canonical vectors have singularities at the origin and require high-degree polynomials to capture the behavior. In Fig. 2 right, we plot the Chebyshev approximation error of different canonical vectors for a fixed Chebyshev degree. We can see given a fixed degree $m=129$, the error of the short-range functions is much larger than for the long-range functions with a jump around $k=15$. The same behavior can also be seen from Fig. 3, in which the absolute values of the Chebyshev coefficients of the interpolant of a long-range vector ${\bf p}_{10}$ and a short-range vector ${\bf p}_{30}$ are shown. We can see that for the long-range vector, the Chebyshev coefficients decay exponentially while for the short-range vector, the coefficients do not decay very much.
@@
Figure 2: Relative $\ell_{\infty}$ error Eq. 37 of the Chebyshev interpolation of the canonical vectors vs. the degree of the Chebyshev interpolant. Each line represents the error of the Chebyshev interpolant of a canonical vector ${\bf p}_{k}$ for $k=1,\ldots,15$ (left) and $k=16,\ldots,39$ (middle). Right: Relative $\ell_{\infty}$ error of the degree $m=129$ Chebyshev interpolation of the canonical vectors vs. the index $k$ of the univariate function.
Figure 3: Absolute values of the Chebyshev coefficients of the degree-129 Chebyshev interpolant of ${\bf p}_{10}$ (left) and ${\bf p}_{30}$ (right). Only coefficients of even degree are shown, i.e., ${\bf c}_{k}(0),{\bf c}_{k}(2),\ldots,{\bf c}_{k}(128)$ since $p(\|x\|)$ is even and ${\bf c}_{k}(1),\ldots,{\bf c}_{k}(129)=0$
@@ 4.1.3 Range-separation of the CP tensor ${\bf P}_{R}$
For a given appropriately chosen range-separation parameter $R_{s}\in[1,R]\cap\mathbb{Z}$ (see [6] for details of the strategies for choosing $R_{s}$), we further use the range-separated representation of the CP tensor ${\bf P}_{R}$ constructed in [6] in the form ${\bf P}_{R}={\bf P}_{R_{s}}+{\bf P}_{R_{l}}$, where
(38) ${\bf P}_{R_{s}}={\sum}_{k=1}^{R_{s}}{\bf p}_{k}\otimes{\bf p}_{k}\otimes{\bf p}_{k},\ {\bf P}_{R_{l}}={\sum}_{k=R_{s}+1}^{R}{\bf p}_{k}\otimes{\bf p}_{k}\otimes{\bf p}_{k}.$
Here, Eq. 38 represents the sums over the sets of indexes for the long- and short-range canonical vectors in the range-separation splitting of the Newton kernel. Then the short-range part ${\bf P}_{R_{s}}$ is a highly localized tensor which has very small support and can thus be stored sparsely (see [6] for details), while the ChebTuck approximation can be gainfully applied to the long-range part ${\bf P}_{R_{l}}$.
We apply Algorithm 3 to the long-range part ${\bf P}_{R_{l}}$ and measure the errors of the approximation at the middle slice as in Eq. 36 for different grid size $n$ and Chebyshev degree $m$, which are shown in Table 2. As opposed to Table 1, we can see that the ChebTuck approximation of the long-range part ${\bf P}_{R_{l}}$ can give satisfactory results already for reasonable Chebyshev degrees $m$. We also show the optimal Tucker ranks, i.e., size of the core tensor $\bm{\beta}$ of the ChebTuck approximation of the long-range part ${\bf P}_{R_{l}}$ computed by RHOSVD with a relative error of $10^{-7}$ for different grid size $n$ and Chebyshev degree $m$ in the last column of Table 2. For fixed Chebyshev degree, the Tucker rank of the long-range part only increases logarithmically with the grid size $n$. This indicates that the long-range part can be well approximated by a ChebTuck format with a small number of parameters.
As a side note, we remark that the numerical evidence in Sections 4.1.2 and 4.1.3 aligns with the intuition underlying the fast multipole method (FMM) [26]: the far-field (long-range) component of the potential admits an accurate approximation by low-degree polynomials.
$n$ $m$ 129 257 513 1025 2049 4097 Tucker rank
256 8.4e-8 7.4e-8 7.3e-8 4.9e-9 1.6e-9 1.8e-10 9
512 5.2e-8 5.2e-8 4.6e-8 4.6e-8 3.0e-9 9.9e-10 11
1024 9.4e-9 9.4e-9 9.4e-9 8.7e-9 8.6e-9 5.7e-10 12
2048 3.6e-9 1.7e-9 1.7e-9 1.7e-9 1.6e-9 1.6e-9 13
4096 1.1e-4 7.5e-10 7.4e-10 7.4e-10 7.4e-10 7.1e-10 15
Table 2: Relative $\ell_{\infty}$ error as in Eq. 36 and in the middle slice and the Tucker rank $\max_{\ell=1,2,3}{r_{i_{\ell}}}$ of the ChebTuck approximation of the long-range part ${\bf P}_{R_{l}}$ for different grid sizes $n$ and degree $m$. The Tucker rank is the same for different $m$ for a fixed grid size $n$, therefore only one value is shown.